Problem 4 A sample of 15 small bags of the same brand of candies was selected. Assume that the population distribution of bag weights is normal. The weight of each bag was then recorded. The mean weight was 2 ounces with a sample standard deviations 0.2 ounces. We would like to calculate an 80% confidence interval for the average weight of a sample of size 15 . a. (3%) standard error- b. (396) The critical t value for an 80% confidence interval is erit С. (396) EBM d. (4%) An 80% confidence interval for the population average weight of the candles is from to
Answer #1

Solution:-

a) S.E = 0.05164

b) The critical value for an 80% confidence interval is + 1.345.

c) EBM = + 0.0695

EBM = tcritical*S.E

d) 80% confidence interval for the population average weight of the candies is from C.I = ( 1.9305, 2.0695).

C.I = \bar{x}\pm t_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 2 + 1.345*0.05164

C.I = 2 + 0.0695

C.I = ( 1.9305, 2.0695)